A die is rolled 16 times.
Find the probability of getting at least one 6. Please round the final answer to 2 or 3 decimal places.

Answer by stanbon(75887)

A die is rolled 16 times.
Find the probability of getting at least one 6. Please round the final answer to 2 or 3 decimal places.
-------
Binomial Problem with n = 16 and p(6) = 1/6 ; p(not 6) = 5/6
----------------
P(at least 1 six) = 1 - P(not 6) = 1-(5/6)^16 = 0.9459..
--------------------
Cheers,
Stan H.
------